what is fhe combined mass in grams of objects that have masses of 0.2027 kg, 80.0 g,524 mg?
Answer #1

I kg = 1000gm. 0. 2027 kg = 10. 2027 X1000) g = 202-7 gm. a. Img = 10-3 g . 524 ang = 524 x 103 g = 0.5249. Combined mass = (hope you will like my answer